What is the concentration of M2+ in a saturated solution of a metal hydroxide M(OH)2 (assume Ksp = 1.8 x 10-16) in an aqueous buffer of pH = 8.3?

-The answer I got was 4.5e-5 M. Just wanted to confirm this was correct, thank you!
Answer #1

Given : pH = 8.3

pOH = 14 - pH

pOH = 14 - 8.3

pOH = 5.7

[OH-] = 10-pOH

[OH-] = 10-5.7

[OH-] = 2.0 x 10-6 M

Ksp M(OH)2 = [M+][OH-]2

Substituting the values of Ksp and [OH-]

1.8 x 10-16 = [M+] * (2.0 x 10-6 M)2

[M+] = (1.8 x 10-16) / (2.0 x 10-6 M)2

[M+] = 4.5 x 10-5 M